# Model averaging with mixed criteria for estimating high quantiles of extreme values: Application to heavy rainfall

### v1

@@ 4.2 Simulation setting
To evaluate the performance of the MA methodology, we conducted a simulation study in which high quantiles were known by generating random numbers from a GEVD. This study estimates the 100-year and 200-year return levels (corresponding to the 0.99 and 0.995 quantiles) using a sample size of $n=50$. The number of submodels ($K$) was set to 12, and the number of bootstrap samples was 500. We generated 1,000 random samples to compute the following evaluation measures for comparing different estimators:
@@
The upper endpoint of the GEVD is bounded when $\xi>0$ (in Hosking-Wallis notation). As a result, estimating high quantiles of the GEVD when $\xi>0$ is relatively straightforward, and the RMSE values tend to be smaller. The MLE and LME typically perform well when $\xi\geq 0$. Therefore, we focus on cases when $\xi$ ranges in $(-0.5,0.0)$. The case $\xi\leq-0.5$ is not considered because the GEVD has no finite variance in this range. We conduct experiments for 10 values of $\xi$: $\xi=-0.45,\ -0.4,\ -0.35,\ -0.3,\ -0.25,\ -0.2,\ -0.15,\ -0.1,\ -0.05$, $-0.001$. For better interpretability, we set $\mu=100$ and $\sigma=30$, as estimates of these parameters are equivariant (Casella & Berger 2001). Setting $\mu=0$ and $\sigma=1$ often results in values less than one with decimal points, making distinguishing differences among methods in the summary table harder. Additionally, this (0,1) scaling occasionally produces negative values, which in rare cases can cause numerical issues in computing the likelihood function or inverting the Hessian matrix.
@@ 4.3 Simulation results
Table 2: Simulation results of the bias, standard error (SE), and root mean squared error (RMSE) of the GEVD, computed for the quantile $q_{0.99}$, using the 10 estimation methods across 10 values of $\xi$. The sample size is $n=50$, with parameters $\mu=100$ and $\sigma=30$. Acronyms of the method names are provided in Table 1.
Measure Method .$-.45$$-.4$$-.35$$-.3$$-.25$$-.2$$-.15$$-.1$$-.05$ 00
@@
Table 2 and Figure 2 present the simulation results for bias, SE, and RMSE computed for the quantile $q_{0.99}$ using 10 estimation methods. Acronyms for the method names are provided in Table 1. Since the results for the ‘gLd’ and ‘med’ weighting schemes were similar in this simulation study, only the results for the ‘gLd’ weight are included in Table 2.
@@
Regarding bias, the MLE exhibits overestimation (positive bias), particularly when $\xi\leq-0.15$, while the proposed MA methods with ‘gLd’ weight and MLE.CD display underestimation (negative bias), especially for $\xi\leq-0.3$. The LME, Re.MLEs, and MA.likes generally perform well. For cases where $-0.25<\xi<0$ (moderately heavy-tailed distributions), the LME, MA.gLds, and MA.like0 perform well, producing small biases. For $-0.35\leq\xi\leq-0.25$, MA.like0, Re.MLE2, and LME show good performance. For extremely heavy-tailed case ($\xi<-0.35$), Re.MLE1 and MA.like1 perform well in terms of bias.
Regarding SE and RMSE, for $\xi\leq-0.25$, MA.gLds and MLE.CD exhibits smaller SE values, resulting in lower RMSE than other methods. However, this reduction in SEs is primarily due to severe underestimation, making the lower RMSE in these cases less meaningful. Nevertheless, it is noteworthy that MA.gLds remains both the smallest SE and bias for $-0.25<\xi\leq 0$, leading to genuinely lower and meaningful RMSEs. For $-0.35\leq\xi\leq-0.25$, MA.likes, Re.MLE2, and LME yield smaller RMSE values than other methods. Determining the best-performing methods for $\xi<-0.35$ is difficult, but LME, Re.MLEs, and MA.likes are acceptable choices.
Based on these findings, the most suitable method depends on the value of $\xi$: MA.gLds for $-0.25<\xi\leq 0$, MA.likes, Re.MLE2, and LME for $-0.35\leq\xi\leq-0.25$, and MA.like1, Re.MLE1, and LME for $\xi<-0.35$. Considering bias and RMSE, MLE and MA.cvt do not perform well for the range over $\xi<0$. Without specifying the $\xi$ value, Re.MLEs, LME, and MA.likes are generally reliable methods.
@@ 5 Real data application
For a real-data example, we analyzed a time series of annual maximum daily precipitation (unit: $mm$) recorded in Hae-nam, Korea. The dataset consists of 52 observations from 1971 to 2022 and is available at https://github.com/yire-shin/MA-gev/tree/main/data. We selected this weather station because the data exhibits a heavy right tail, making it well-suited for evaluating the effectiveness of the proposed methods. Figure 3 displays the relative frequency histogram with density curves and a scatterplot with lines representing the 50-year return level estimates.
@@
Figure 3: (Left) Density plots overlaid on the relative frequency histogram and (Right) a scatterplot of observations by year with horizontal lines representing the 50-year return level estimates. The data depicts the annual maximum daily rainfall (unit: $mm$) in Hae-nam, Korea. Both figures’ blue solid line and red dashed line correspond to the estimates obtained using the LME and MLE methods.
Table 3 presents the parameter estimates, 100-year return levels, and their SE obtained using the formula (21) and nonparametric bootstrap (NpB) with 500 resamples for 10 estimation methods. Parameter estimates for the MA methods are derived from the surrogate models. In Table 3, the asymptotic SEs of MA.gLds (MA.likes) obtained using the formula (21) are greater (less) than the corresponding SEs computed via nonparametric bootstrap. The SE (NpB) of MA methods is smaller than other methods, whereas MLE, MLE.CD, and Re.MLE1 yield larger SEs. The SEs of Re.MLE2 and LME fall within the moderate range—the 100-year return levels estimated by MA.likes and Re.MLE2 lies between those obtained using LME and MLE. Based on the findings from the simulation study, the return levels of MA.gLds may be underestimated despite their small SE (NpB) values. Conversely, the return level estimated by MLE may be overestimated, with exceptionally large SEs in both asymptotic and bootstrap calculations.
@@
Figure 4 presents the sampling distributions of 100-year return levels from nine estimation methods, displayed as density plots over the relative frequency histogram, based on 1,000 bootstrap samples. Except for MLE, all distributions appear approximately normal, though some exhibit slight right skewness.
Figure 4: Sampling distributions of 100-year return levels from nine estimation methods, displayed as density plots over the relative frequency histogram. These distributions were obtained using 1,000 bootstrap samples from the annual maximum daily rainfall data (unit: $mm$) in Hae-nam, Korea. Acronyms for the method names are provided in Table 1.
Figure 5 presents quantile-per-quantile plots for nine estimation methods. We used the standard plotting position formula: $q_{i}=(i-0.5)/n$ to construct these plots. Surrogate models were utilized for the MA methods. The plots for MLE and Re.MLE1 appear to overestimate the largest observation, whereas the other methods fit the data relatively well.
@@
6 Discussion
6.1 Bias correction and Bayesian model averaging
The proposed MA methods with ‘gLd’ and ‘med’ weights exhibit negative biases despite the left-trimming approach adopted in this study, mainly when $\xi$ is less than $-0.3$. To address this issue, a bias correction technique may be necessary. Bayesian model averaging (BMA) could provide a solution for correcting such biases. BMA assigns weights proportional to the model’s marginal likelihood (Hoeting et al. 1999; Raftery et al. 2005; Shin et al. 2019; Vettori et al. 2020, for example among many). The BMA framework begins by specifying prior probabilities $p(M_{k})$ for all candidate models $M_{1},\dots,M_{K}$ and defining the prior density $p(\>\underline{x}|M_{k})$. The distribution of the return level given the data ($\>\underline{x}$) is then expressed as:
@@ 6.1 Bias correction and Bayesian model averaging
where $p(\>\underline{x}|M_{k})$ is typically the likelihood of model $M_{k}$ and $p(M_{k})$ represents the prior probability of $M_{k}$. The posterior mean and variance of the BMA-predicted return level are given by (Hoeting et al. 1999; Claeskens & Hjort 2008)
@@
where weight $w_{k}$ is the same to (30). For estimating return level in the GEV, we estimate $E(r|M_{k},\>\underline{x})$ by $r(\hat{\mu}_{2}(\xi_{k}),\hat{\sigma}_{2}(\xi_{k}),\xi_{k})$ in this study. Then the estimator of $E(r|\>\underline{x})$ in (31) is the same as $\hat{r}_{\text{MA}}$ in (8), but with different weights. The weights in (7) and (24) do not use the prior $p(M_{k})$, which is different from the weights in (31).
For the prior of $M_{k}$, we consider the prior for the shape parameter $\xi$ only, which is similar to a reversed form of the penality function of Coles and Dixon (1999):
$p(M_{k})\ =\ p(\xi_{k})\ =\ \begin{cases}\ 1&~{}\mbox{ if~{} $\xi_{k}\leq-1\ $or $\ \xi_{k}\geq 0$}\\ \ 1-2\xi_{k}&~{}\mbox{ if~{} $-1.0<\xi_{k}<0$}.\end{cases}$ (32)
We employed this function mainly for bias correction of $\hat{r}_{\text{MA}}$ because our estimator $\hat{r}_{\text{MA}}$ with ‘gLd’ weighting scheme turned out to underestimate the true return level for negative $\xi$ (especially for $\xi\leq-0.3$) in the simulation study. By providing more weight for $\xi_{k}$ closer to $-1.0$, we expected the return level estimator to correct the underestimation bias of $\hat{r}_{\text{MA}}$.
The results of a simulation study using BMA with the above prior (32) showed only a little (about 5-7%) reduction of biases for $\xi\leq-0.3$ and the new addition of overestimation biases for $\xi>-0.3$. In other words, we did not successfully correct the underestimation bias of $\hat{r}_{\text{MA}}$ using the above BMA scheme, so we do not document the results in this report. However, we believe that a more elaborate application of BMA than that employed in this study can find an answer to improve the estimation of high quantiles with accuracy.

### v2

@@ 4.2 Simulation setting
To evaluate the performance of the MA methodology, we conducted a simulation study by generating random numbers from a GEVD in which high quantiles were known. This study estimated the 100-year and 200-year return levels (corresponding to the 0.99 and 0.995 quantiles) using a sample size of $n=50$. The number of submodels ($K$) was set to 12, and the number of bootstrap samples was 500. We generated 1,000 random samples to compute the following evaluation measures for comparing different estimators:
@@
The upper endpoint of the GEVD is bounded when $\xi>0$ (under the Hosking-Wallis notation). As a result, estimating high quantiles of the GEVD when $\xi>0$ is relatively straightforward, and the RMSE values tend to be smaller. The MLE and LME typically perform well when $\xi\geq 0$. Therefore, we focus on cases when $\xi$ ranges over (0.5,0.0)(-0.5,0.0). The case $\xi\leq-0.5$ is not considered because the GEVD has no finite variance in this range. We conduct experiments for 10 values of $\xi$: $\xi=-0.45,\ -0.4,\ -0.35,\ -0.3,\ -0.25,\ -0.2,\ -0.15,\ -0.1,\ -0.05$, and 0.001-0.001. For better interpretability, we set $\mu=100$ and $\sigma=30$, as estimates of these parameters are equivariant (Casella & Berger 2001). Setting $\mu=0$ and $\sigma=1$ often results in values less than one with decimal points, making it harder to distinguish differences among methods in the summary table. Additionally, this (0,1) scaling occasionally produces negative values, which in rare cases can cause numerical issues in computing the likelihood function or inverting the Hessian matrix.
@@ 4.3 Simulation results
Table 2: Simulation results of the bias, standard error (SE), and root mean squared error (RMSE) of the GEVD, computed for the quantile $q_{0.99}$, using the ten estimation methods across ten values of $\xi$. The sample size is $n=50$, with parameters $\mu=100$ and $\sigma=30$. Acronyms of the method names are provided in Table 1.
Measure Method .45-.45 .4-.4 .35-.35 .3-.3 .25-.25 .2-.2 .15-.15 .1-.1 .05-.05 0
@@
Table 2 and Figure 2 present the simulation results for bias, SE, and RMSE computed for the quantile $q_{0.99}$ using ten estimation methods. Acronyms for the method names are provided in Table 1. Since the results for the ‘gLd’ and ‘med’ weighting schemes were similar in this simulation study, only the results for the ‘gLd’ weighting scheme are included in Table 2.
@@
Regarding bias, the MLE exhibits overestimation (positive bias), particularly when $\xi\leq-0.15$, while the proposed MA methods with ‘gLd’ weight and MLE.CD display underestimation (negative bias), especially for $\xi\leq-0.3$. The LME, Re.MLEs, and MA.likes generally perform well. For cases where $0.25<\xi<0$ (moderately heavy-tailed distributions), the LME, MA.gLds, and MA.like0 perform well, producing small biases. For $0.35\leq\xi\leq-0.25$, MA.like0, Re.MLE2, and LME show good performance. For extremely heavy-tailed cases ($\xi<-0.35$), Re.MLE1 and MA.like1 perform well in terms of bias.
Regarding SE and RMSE, for $\xi\leq-0.25$, MA.gLds and MLE.CD exhibit smaller SE values, resulting in lower RMSE values than other methods. However, this reduction in SEs is primarily due to severe underestimation, making the lower RMSE in these cases less meaningful. Nevertheless, it is noteworthy that MA.gLds remains both smallest in SE and bias for $0.25<\xi\leq 0$, leading to genuinely lower and meaningful RMSEs. For $0.35\leq\xi\leq-0.25$, MA.likes, Re.MLE2, and LME yield smaller RMSE values than other methods. Determining the best-performing methods for $\xi<-0.35$ is difficult, but LME, Re.MLEs, and MA.likes are acceptable choices.
Based on these findings, the most suitable method depends on the value of $\xi$: MA.gLds for $0.25<\xi\leq 0$, MA.likes, Re.MLE2, and LME for $0.35\leq\xi\leq-0.25$, and MA.like1, Re.MLE1, and LME for $\xi<-0.35$. Considering bias and RMSE, MLE and MA.cvt do not perform well for the range of $\xi<0$. Without specifying the $\xi$ value, Re.MLEs, LME, and MA.likes are generally reliable methods.
@@ 5 Real data application
For a real-data example, we analyzed a time series of annual maximum daily precipitation (unit: $mm$) recorded in Hae-nam, Korea. The dataset consists of 52 observations from 1971 to 2022 and is available at the GitHub repository https://github.com/yire-shin/MA-gev/tree/main/data. We selected this weather station because the data exhibit a heavy right tail, making it well-suited for evaluating the effectiveness of the proposed methods. Figure 3 displays the relative frequency histogram with density curves and a scatterplot with lines representing the 50-year return level estimates together with shaded regions for 90% confidence intervals.
@@
Figure 3: (Left) Density plots overlaid on the relative frequency histogram and (Right) a scatterplot of observations by year with horizontal lines representing the 50-year return level estimates together with shaded regions for 90% confidence intervals. The data depict the annual maximum daily rainfall (unit: $mm$) in the Hae-nam station, Korea. In both figures, the blue solid line and the red dashed line correspond to the estimates obtained using the LME and MLE methods, respectively.
Table 3 presents the parameter estimates, 100-year return levels, and their SEs obtained using the formula (22) and the nonparametric bootstrap (NpB) with 500 resamples for 10 estimation methods. The number of submodels ($K$) was set to 12. Parameter estimates for the MA methods are derived from the surrogate models. In Table 3, the asymptotic SEs of MA.gLds (MA.likes) obtained using the formula (22) are greater (less) than the corresponding SEs computed via nonparametric bootstrap. The SE (NpB) of MA methods is smaller than that of other methods, whereas MLE, MLE.CD, and Re.MLE1 yield larger SEs. The SEs of Re.MLE2 and LME fall within the moderate range—the 100-year return levels estimated by MA.likes and Re.MLE2 lie between those obtained using LME and MLE. Based on the findings from the simulation study, the return levels of MA.gLds may be underestimated despite their small SE (NpB) values. Conversely, the return level estimated by MLE may be overestimated, with exceptionally large SEs in both asymptotic and bootstrap calculations.
@@
Figure 4 presents the sampling distributions of the 100-year return levels from nine estimation methods, displayed as density plots overlaid on the relative frequency histogram, based on 1,000 bootstrap samples. Except for the MLE, all distributions appear approximately normal, though some exhibit slight right skewness.
Figure 4: Sampling distributions of the 100-year return levels from nine estimation methods, displayed as density plots overlaid on the relative frequency histogram. These distributions were obtained using 1,000 bootstrap samples from the annual maximum daily rainfall data (unit: $mm$) in Hae-nam, Korea. Acronyms for the method names are provided in Table 1.
Figure 5 presents quantile-per-quantile plots for nine estimation methods. We used the standard plotting position formula: $q_{i}=(i-0.5)/n$ to construct these plots. Surrogate models were utilized for the MA methods. The plots forthe MLE and Re.MLE1 appear to overestimate the largest observation, whereas the other methods fit the data relatively well.
@@
6 Additional study
6.1 Selection of the number of submodels ($K$)
In this section, we propose a systematic way to select an appropriate $K$. If $K$ is too large, unnecessary computation is required; if $K$ is too small, the variation of MA estimators due to the use of different $K$ values can be high. A sensitivity analysis by varying $K$ from 4 to 20 provides information on the performance of the estimation method with $K$. We propose a method considering two measures simultaneously. The first measure is the stability of the (100-year) return level estimates as $K$ varies from 4 to 20. The second measure is the standard error of the return level estimate.
We define the stability measure of the return level estimate with respect to $K$ as
$d_{K}=|\hat{r}_{\text{MA}}(K)\,-\,\hat{r}_{\text{MA}}(K-1)|+|\hat{r}_{\text{MA}}(K+1)\,-\,\hat{r}_{\text{MA}}(K)|,$ (30)
for $K=5,\dots,19$. The smaller $d_{K}$ means that $\hat{r}_{\text{MA}}(K)$ is more stable. Figure 7 shows a sensitivity analysis by varying $K$ from 4 to 20, with the MA.like1 method for the Hae-nam data. The black dots indicate the 100-year $\hat{r}_{\text{MA}}(K)$. The red solid line with triangular markers indicates the stability measures $d_{K}$ rescaled to the range of $\hat{r}_{\text{MA}}(K)$. We treat the estimate as stable when $d_{K}$ is less than or equal to the $\alpha$-quantile of $d_{K},\,K=5,\dots,19$. We employed $\alpha=0.6$ in this study. The $\alpha$-quantile is denoted by $q_{\alpha}(d_{K})$. The orange horizontal dashed line stands for $q_{.6}(d_{K})$ of rescaled $d_{K}$ values. In Figure 7, the estimates for $K\leq 10$ are unstable in terms of $d_{K}$. Whereas, the estimates for $K\geq 11$ are stable. We will choose an appropriate $K$ among these values which lead to stable estimates. Let us denote a set of such $K$s by $I_{S}=\{K:d_{K}\leq q_{\alpha}(d_{K})\}$.
Figure 7: A sensitivity analysis to select the optimal $K$, applied to the annual maximum daily rainfall data (unit: $mm$) from Hae-nam, Korea. The black dots indicate the 100-year return level estimates obtained by the MA.like1 method, the red solid line with triangular markers indicates the stability measures defined in (30), and the blue solid line with circular markers depicts the standard errors of return level estimates.
For the second measure, we want to select the $K$ such that the standard error (SE) of $\hat{r}_{\text{MA}}(K)$ is relatively small. The square root of the formula (24) is employed for computing the SE, which is denoted by $SE_{K}$. We will choose an appropriate $K$ for which $SE_{K}$ is less than or equal to the $\alpha$-quantile of these SEs. We employed $\alpha=0.6$ in this study. The $\alpha$-quantile is denoted by $q_{\alpha}(SE_{K})$. The blue solid line with circular markers in Figure 7 depicts the $SE_{K}$ values rescaled to the range of $\hat{r}_{\text{MA}}(K)$. The green dashed horizontal line stands for $q_{.6}(SE_{K})$ of rescaled $SE_{K}$ values.
Let $I_{E}=\{K:SE_{K}\leq q_{\alpha}(SE_{K})\}$. The $K^{\prime}$ is selected as the minimum among those $K$s that satisfy the stability and SE conditions simultaneously. In other words,
$K^{\prime}=\text{min}\,\{K:\,K\in I_{S}\cap I_{E}\}.$ (31)
Then the optimal $K^{*}$ is selected as the $K$ which leads to a smaller standard error near $K^{\prime}$, for example, among $K^{\prime}_{N}=(K^{\prime}-1,\,K^{\prime}-2,\,K^{\prime},\,K^{\prime}+1,\,K^{\prime}+2$) which satisfy the stability condition. That is,
$K^{*}=\text{arg}\,\text{min}\,\{SE_{K}\}~~\text{for}~~K\in K^{\prime}_{N}\cap I_{S}.$ (32)
In Figure 7, $K^{\prime}=11$ which coincides with $K^{*}$. The vertical line represents the optimal $K^{*}$. It is notable that $K=17$ has the minimum standard error among stable estimates, but it may be too large and the estimates for $K=11$ and $K=17$ look similar. Thus we chose $K^{*}=11$ as the optimal number of submodels. Another example of choosing the optimal $K$ in this way is presented in the Supplementary Material.
6.2 Bias reduction and Bayesian model averaging
The proposed MA methods with ‘gLd’ and ‘med’ weights exhibit negative biases despite the left-trimming approach adopted in this study, mainly when $\xi$ is less than 0.3-0.3. To address this issue, a bias correction technique may be necessary. Bayesian model averaging (BMA) could provide a solution for correcting such biases. BMA assigns weights proportional to the model’s marginal likelihood (Hoeting et al. 1999; Raftery et al. 2005; Shin et al. 2019; Vettori et al. 2020, for example among many others). The BMA framework begins by specifying prior probabilities $p(M_{k})$ for all candidate models $M_{1},\dots,M_{K}$ and defining the prior density $p(\>\underline{x}|M_{k})$. The distribution of the return level given the data ($\underline{x}$) is then expressed as:
@@ 6.2 Bias reduction and Bayesian model averaging
where $p(\>\underline{x}|M_{k})$ is typically the likelihood of model $M_{k}$, which corresponds to (6) for MA.gLd method and $p(M_{k})$ represents the prior probability of $M_{k}$. The posterior mean and variance of the BMA-predicted return level are given by (Hoeting et al. 1999; Claeskens & Hjort 2008)
@@
where the weight $w_{k}$ is the same as (34). For estimating the return level in the GEV, we estimate $E(r|M_{k},\>\underline{x})$ by $r(\hat{\mu}(\xi_{k}),\hat{\sigma}(\xi_{k}),\xi_{k})$ in this study. Then the estimator of $E(r|\>\underline{x})$ in (35) is the same as $\hat{r}_{\text{MA}}$ in (8), but with different weights. The weights in (7) and (25) do not use the prior $p(M_{k})$, which is different from the weights in (35). The first term on the right hand side of variance formula (35) is the among-model variance, and the second term is the within-model variance. In this study, $Var(r|\>\underline{x},M_{k})$ is calculated using the asymptotic variance formula (13).
The selection of a prior is important to reduce the bias. We consider the prior for the shape parameter $\xi$ only, not for the location and scale parameters, because $\xi$ is unit-free and affects return level estimation sensitively. Indeed, some authors such as Coles and Dixon (1999), Martins and Stedinger (2000), and Cannon (2010) considered a prior or a penalty function for $\xi$ only. Whereas, Lee et al. (2017) and Shin et al. (2025b) suggested an approach using the data-adaptive penalty function (PF) which provides different hyperparameters according to an earlier estimate (LME in this study) of $\xi$. The approach using the data-adaptive PF turned out to be effective in correcting the estimation bias and robust to a poor selection of the prior (Lee et al. (2017); Shin et al. (2025b)). The data-adaptive PF may be viewed as an empirical prior.
We considered a PF based on the normal density:
$p(M_{k})\ =\ p(\xi_{k})\ =\psi(\xi_{k};\>\mu_{\xi},\sigma_{\xi})$ (36)
where $\psi(\xi_{k};\>\mu_{\xi},\sigma_{\xi})$ represents a normal density of $\xi_{k}$ with a mean $\mu_{\xi}$ and standard deviation $\sigma_{\xi}$. The behavior of this PF depends on the choice of hyperparameters, $\mu_{\xi}$ and $\sigma_{\xi}$. We chose the following hyperparameters, varying by the weighting method: For MA.gLd or MA.med,
[…]
where $\hat{\xi}$ is the LME of $\xi$. For the MA.like method,
[…]
The above PF is dependent on $\hat{\xi}$ (LME of $\xi$), because the underestimation bias of MA estimate is larger (smaller) for more (less) negative values of $\xi$. By providing more weight for $\xi_{k}$ closer to 1.0-1.0, we expect the MA method to correct the underestimation bias of $\hat{r}_{\text{MA}}$. When $\hat{\xi}\geq 0$, we set $p(\xi_{k})=1/K$ for all $k$. Some (five) constants in the above PFs were acquired by tuning the PF so as to approximately minimize the biases, using Monte Carlo simulations several times. One can of course use different tuning constants or different PFs. See Shin et al. (2025b) for a data-adaptive beta PF.
Figure 8 shows graphs of these PFs for weighting methods (MA.like and MA.gLd) and for different estimates of $\xi$ such as 0.1,0.2,0.3-0.1,\,-0.2,\,-0.3, and 0.4-0.4, respectively. As $\hat{\xi}$ decreases, a more pronounced shift to the left is assigned, so as to differentially correct the underestimation biases according to the value of $\hat{\xi}$.
Figure 8: Graphs of penalty functions which are different for weighting methods (like and gLd) and for $\hat{\xi}$ (LME). The vertical dashed line represents the LME obtained from data which are 0.1,0.2,0.3-0.1,\,-0.2,\,-0.3, and 0.4-0.4, respectively.
Table 4 shows the results of a simulation study using BMA with the above PFs, acquired under the same settings as in Section 4. Compared to Table 2, considerable reduction of underestimation biases is achieved for $\xi\leq-0.3$. However, new overestimation biases are observed for $\xi\geq-0.25$. The SE and RMSE of BMA.like are reduced from those of MA.like, and are comparable with or smaller than those of Re.MLE1 and LME. This result (the reductions of bias and SE) is a remarkable advantage of using BMA.like. Whereas, the SE and RMSE of BMA.gLd are increased relative to those of MA.gLd.
Table 4: Simulation results same as Table 2 but for Bayesian model averaging methods.
Measure Weight -0.45 -0.40 -0.35 -0.3 -0.25 -0.2 -0.15 -0.1 -0.05 0.0
Bias BMA.like -13.4 -2.0 1.2 3.8 11.8 8.5 6.9 1.6 1.8 -1.5
Bias BMA.gLd -16.8 -3.5 -2.2 2.7 11.7 7.9 4.3 -0.3 -0.9 -2.4
SE BMA.like 174.9 156.3 125.8 100.8 99.3 75.5 62.0 47.2 38.3 31.8
SE BMA.gLd 186.6 170.4 142.2 119.2 114.2 83.4 66.7 49.6 37.6 30.0
RMSE BMA.like 175.4 156.3 125.8 100.9 100.0 76.0 62.4 47.2 38.3 31.8
RMSE BMA.gLd 187.4 170.4 142.2 119.2 114.8 83.8 66.8 49.6 37.6 30.1
Table 5 provides the estimates and SEs calculated by BMA methods for Hae-nam data. Compared to Table 3, the 100-year return levels are increased from MA methods. The ‘SE among model’ and ‘SE within model’ are calculated by taking the square roots of the first and second terms of the right hand side of the variance formula (35), respectively. The ‘Total SE’ is obtained by taking the square root of the variance formula (35). The (Total) SE of BMA.gLD increased from that of MA.gLd, whereas the SE of BMA.like is decreased from that of MA.like. In BMA.like, the among-model variance is relatively small compared to the within-model variance, which results in a decrease of total SE. In contrast, the among-model variance in BMA.gLd is relatively large, which results in an increase of total SE.
Although the BMA methods considered here showed a substantial improvement, we believe that more elaborate choices of PF and tuning constants could lead to more accurate estimation of return levels than the BMA method presented here.
Table 5: Same as Table 3 but for Bayesian model averaging methods. The ‘SE am(ong).model’ and ‘SE wi(thin).model’ are calculated by taking the square roots of the first and second terms of the right hand side of the variance formula (35), respectively. The ‘Total SE’ is obtained by taking the square root of the variance formula (35).
Method $\hat{\mu}$ $\hat{\sigma}$ $\hat{\xi}$ 100-y RL SE am.model SE wi.model Total SE
BMA.like 113.59 34.87 -0.348 520.59 14.73 60.29 62.07
BMA.gLd 116.43 34.02 -0.321 507.70 65.34 58.66 87.81
6.3 Guide for using MA methods
  1. 1.
    Starter method: MLE. The starter LME depends on the confidence interval obtained from the bootstrap, which may vary slightly across different runs. Whereas, the starter MLE rarely depends on this randomness.
  2. 2.
    Selection of $K$: Employ the method described in Section 6.1.
  3. 3.
    Weighting criterion: MA.like1 or BMA.like. In general, the MA.like1 seems to perform well in the simulation study. If the user is uncomfortable with trimming the lowest observation or prefers a Bayesian approach, the BMA.like method can be a good alternative.
7 Discussion
7.1 Bias correction
We considered the reasons why the MA methods proposed in this study underestimate the return level. One plausible answer can be found in the fact that the LME sometimes produces underestimation (UE) bias for some negative $\xi$ (especially for $\xi\leq-0.3$). Consequently, the MA.like method results in UE, because it may accumulate UE by computing the LME $K$ times. In the MA.gLd method, the covariance matrix $V$ was used in constructing $\text{GLD}_{k}$ in (5) over all $k$, instead of using $V_{k}$. This uniform usage might cause a poor approximation to $\text{GLD}_{k}$, especially for large negative $\xi_{k}$. Further investigation is needed for this undesirable bias.
In addition, we employed the multivariate normal density (6) to approximate the likelihood function of population L-moments given sample L-moments. This approximation may be valid for a large sample. But, it may not perform well for a small sample, which probably induces UE bias by assigning too small $p(\>\underline{x}|M_{k})$ values for large negative $\xi_{k}$. We may need to use an asymmetric distribution or the multivariate $t$-distribution instead of the normal density (6).